# Phonon effects, impact ionization and power conversion in Mott photovoltaic systems

### v1

@@ Appendix B The Hartree shift for acoustic phonons
The starting point for the derivation of the Hartree shift $\varepsilon_{\textrm{H},z}$ per spin per layer is the expression for the singular part of $\Sigma_{\text{e-ph},z}^{\text{R}}(z)$, in which $z$ as subscript indicates the layer index and $z$ in the round brackets the Keldysh contour time argument of the SE. This quantity reads
@@
where in the last line we use that $G_{\text{loc},zz}^{<}(t^{\prime};{t^{\prime}})=\operatorname*{i}n_{z}$ at the NESS. The starting point $t_{0}$ of the Keldysh contour may be taken as $t_{0}=0$ or $t_{0}=-\infty$ with no influence on the final result. Using Eq. (13) for $D_{\text{ph}}^{\textrm{R}}(t,t^{\prime})$ and $t_{0}=0$, after some standard integrations we arrive to
$\Sigma_{\text{e-ph},z,\textrm{H}}^{\textrm{R}}(t)=-\frac{4g^{2}n_{z}}{\omega_{ph}^{3}}\frac{t^{2}\omega_{ph}^{4}}{1+t^{2}\omega_{ph}^{2}}.$ (27)
@@
$\varepsilon_{\textrm{H},z}=\lim_{t\rightarrow\infty}\Sigma_{\text{e-ph},z,\textrm{H}}^{\textrm{R}}(t)=-\frac{2g^{2}}{\omega_{ph}}\Delta n_{z},$ (28)
in which we use Eq. (21). This quantity, which vanishes at half-filling, has to be computed self-consistently in the DMFT loop  [^19] The exceeding charge depends on the number of particles which is, in turn, computed self-consistently at every DMFT iteration..
@@ Appendix C The effective electron-phonon coupling strength
$\alpha^{2}F_{\bf k}(\varepsilon,\omega)=\sum_{\bf q}\delta(\omega-\omega_{\bf q})\sum_{\bf k^{\prime}}|g_{\bf k^{\prime},\bf k}^{\bf q}|^{2}\delta(\varepsilon-\varepsilon_{\bf k^{\prime}}),$ (29)
in which $\bf q$ and $\bf k,\bf k^{\prime}$ are the phonon and electron crystal momenta and $g_{\bf k^{\prime},\bf k}^{\bf q}$ is the general e-ph matrix element. This function provides the information relative to the e-ph coupling of the electronic state $\bf k$. Given the e-ph interaction Hamiltonian in Eq. (3), this expression simplifies to
$\alpha^{2}F(\varepsilon,\omega)=g^{2}\sum_{\bf q}\delta(\omega-\omega_{\bf q})\sum_{\bf k^{\prime}}\delta(\varepsilon-\varepsilon_{\bf k^{\prime}}).$ (30)
We have now the product of the phonon and electron spectral functions separately. For the former we use the acoustic phonons’ DOS $A_{\text{ph}}(\omega)=(\omega/\omega^{2}_{\text{ph}})\mathrm{e}^{-|\omega|/\omega_{\text{ph}}}$ from Sec. III.1, which leads to
$\alpha^{2}F(\varepsilon,\omega)=g^{2}A_{\text{ph}}(\omega)A(\varepsilon)\approx g^{2}\frac{A_{\text{ph}}(\omega)}{W_{\text{b}}},$ (31)
where we approximate the electron DOS $A(\varepsilon)$ with the inverse of the Hubbard bands’ bandwidth $W_{\text{b}}$, introduced in Sec. IV. The strength of e-ph coupling is then encoded in the dimensionless parameter [102]
$\lambda_{\textrm{eff}}=2\int_{0}^{\infty}\frac{d\omega}{\omega}\alpha^{2}F(\omega)=\frac{2g^{2}}{W_{\text{b}}\omega_{\text{ph}}},$ (32)
in which $\alpha^{2}F(\omega)$ from Eq. (31) depends now only on $\omega$. Finally, we notice that the Hartree shift and $\lambda$ are related by:
$\varepsilon_{\textrm{H},z}=-\lambda_{\textrm{eff}}W_{\text{b}}\Delta n_{z}.$ (33)
@@ Appendix D Derivation of the time-averaged absorbed power
In this Appendix we derive the time-averaged absorbed power $\overline{P}_{\text{abs}}$ introduced in Sec. III.2 of the main text.
The time-dependent absorbed power $P_{\text{abs}}(t)$ originates from the oscillating current $\bm{j_{\parallel}}(t)$ along the lattice body diagonal of the $d$-dimensional correlated layers, in the direction of the external electric field $\bm{E}(t)$.
We start from the expression of $P_{\text{abs}}(t)$ as the classical expectation value
$P_{\text{abs}}(t)=\frac{\mathop{}\!\mathrm{d}E_{\text{abs}}}{\mathop{}\!\mathrm{d}t}=\Big{\langle}\frac{\mathop{}\!\mathrm{d}\hat{H}_{\text{H}}(t)}{\mathop{}\!\mathrm{d}t}\Big{\rangle}.$ (34)
The time-dependence in $\hat{H}_{\text{H}}(t)$  [^20] The subscript H stays for the Heisenberg representation. appears only via the vector potential $\bm{A}(t)$ in the intralayer hopping term $t_{z}(t)$ (2), introduced in Sec. II. For this reason, one obtains the general result [63]
$P_{\text{abs}}(t)=\Big{\langle}\frac{\partial\hat{H}_{\text{H}}(t)}{\partial\bm{A}(t)}\cdot\frac{\partial\bm{A}(t)}{\partial t}\Big{\rangle}=\bm{j_{\parallel}}(t)\cdot\bm{E}(t),$ (35)
@@
$\bm{j}_{\parallel}(t)=-\Big{\langle}\frac{\partial\hat{H}_{\text{H}}(t)}{\partial\bm{A}(t)}\Big{\rangle}=\frac{2\operatorname*{i}}{N}\sum_{z=1}^{\text{L}}\sum_{\bm{k}}\bm{v}_{\bm{k}}(t)G_{z,\bm{k}}^{<}(t,t),$ (36)
where $N$ is the number of lattice sites of the single $d$-dimensional correlated layer and $n_{z,\bm{k}}(t)=-\operatorname*{i}G_{z,\bm{k}}^{<}(t,t)$.
@@
$\begin{split}&\overline{P}_{\text{abs}}=\frac{1}{T}\int_{-T/2}^{T/2}\mathop{}\!\mathrm{d}tP_{\text{abs}}(t)\\ &=\frac{2\operatorname*{i}}{N}\sum_{z=1}^{\text{L}}\sum_{\bm{k}}\frac{1}{T}\int_{-T/2}^{T/2}\mathop{}\!\mathrm{d}t\bm{v}_{\bm{k}}(t)\cdot\bm{E}(t)\\ &\sum_{l=-\infty}^{\infty}e^{-\operatorname*{i}l\Omega t}\int_{-\infty}^{+\infty}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}G_{l,z,\bm{k}}^{<}(\omega),\end{split}$ (37)
@@
with $t_{\text{rel}}=t-t^{\prime}$ and $t_{\text{av}}=(t+t^{\prime})/2$ defined in Sec. III.2. The scalar product in (37) reads
$\begin{split}\bm{v}_{\bm{k}}(t)\cdot\bm{E}(t)&=E_{0}\cos(\Omega t)\Big{\{}-\overline{\epsilon}\cos\Big{[}\mathscr{A}\sin(\Omega t)\Big{]}\\ &+\epsilon\sin\Big{[}\mathscr{A}\sin(\Omega t)\Big{]}\Big{\}}\end{split}$ (39)
with the quantities $\mathscr{A}$,$\epsilon$,$\overline{\epsilon}$ defined in Sec. II. After some algebra, one finally obtains for the Wigner and layer resolved integrand in (37)
$\begin{split}\overline{P}_{\text{abs},l,z}(\omega)&=-E_{0}\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\rho(\epsilon,\overline{\epsilon})\Big{(}J_{1-l}(\mathscr{A})+J_{-1-l}(\mathscr{A})\Big{)}\\ &\Big{[}\epsilon(1-\delta_{l,0})\delta_{l,2k}+\operatorname*{i}\overline{\epsilon}\delta_{l,2k+1}\Big{]}G_{l,z}^{<}(\omega;\epsilon,\overline{\epsilon}),\end{split}$ (40)
in which $k\in\mathbb{Z}$, $J_{n}$ is the $n$-th order Bessel function of the first kind and the integration over the JDOS $\rho(\epsilon,\overline{\epsilon})$ is used (see Sec. II), which leads to Eq. (LABEL:eq:avg_abs_pow) in Sec. III.2.

### v2

@@ Appendix B The Hartree shift for acoustic phonons
The starting point for the derivation of the Hartree shift $\varepsilon_{\mathrm{H},z}$ per spin per layer is the expression for the singular part of $\Sigma_{\mathrm{e-ph},z}^{\mathrm{R}}(z)$, in which $z$ as subscript indicates the layer index and $z$ in the round brackets the Keldysh contour time argument of the SE. This quantity reads
@@
[…]
where in the last line we use that $G_{\mathrm{loc},zz}^{<}(t^{\prime};{t^{\prime}})=\mathrm{i}n_{z}$ at the NESS. The starting point $t_{0}$ of the Keldysh contour may be taken as $t_{0}=0$ or $t_{0}=-\infty$ with no influence on the final result. Using Eq. 13 for $D_{\mathrm{ph}}^{\textrm{R}}(t,t^{\prime})$ and $t_{0}=0$, after some standard integrations we arrive to
$\Sigma_{\mathrm{e-ph},z,\mathrm{H}}^{\textrm{R}}(t)=-\frac{4g^{2}n_{z}}{\omega_{ph}^{3}}\frac{t^{2}\omega_{ph}^{4}}{1+t^{2}\omega_{ph}^{2}}.$ (27)
@@
$\varepsilon_{\mathrm{H},z}=\lim_{t\rightarrow\infty}\Sigma_{\mathrm{e-ph},z,\mathrm{H}}^{\textrm{R}}(t)=-\frac{2g^{2}}{\omega_{ph}}\Delta n_{z},$ (28)
in which we use Eq. 21. This quantity, which vanishes at half-filling, has to be computed self-consistently in the DMFT loop  [^19] The exceeding charge depends on the number of particles which is, in turn, computed self-consistently at every DMFT iteration..
@@ Appendix C The effective electron-phonon coupling strength
$\alpha^{2}F_{\mathbf{k}}(\varepsilon,\omega)=\sum_{\mathbf{q}}\delta(\omega-\omega_{\mathbf{q}})\sum_{\mathbf{k}^{\prime}}|g_{\mathbf{k}^{\prime},\mathbf{k}}^{\mathbf{q}}|^{2}\delta(\varepsilon-\varepsilon_{\mathbf{k}^{\prime}}),$ (29)
in which $\mathbf{q}$ and $\mathbf{k},\mathbf{k}^{\prime}$ are the phonon and electron crystal momenta and $g_{\mathbf{k}^{\prime},\mathbf{k}}^{\mathbf{q}}$ is the general e-ph matrix element. This function provides the information relative to the e-ph coupling of the electronic state $\mathbf{k}$. Given the e-ph interaction Hamiltonian in Eq. 3, this expression simplifies to
$\alpha^{2}F(\varepsilon,\omega)=g^{2}\sum_{\mathbf{q}}\delta(\omega-\omega_{\mathbf{q}})\sum_{\mathbf{k}^{\prime}}\delta(\varepsilon-\varepsilon_{\mathbf{k}^{\prime}}).$ (30)
We have now the product of the phonon and electron spectral functions separately. For the former we use the acoustic phonons’ DOS $A_{\mathrm{ph}}(\omega)=(\omega/\omega^{2}_{\mathrm{ph}})\mathrm{e}^{-|\omega|/\omega_{\mathrm{ph}}}$ from Sec. 3.1, which leads to
$\alpha^{2}F(\varepsilon,\omega)=g^{2}A_{\mathrm{ph}}(\omega)A(\varepsilon)\approx g^{2}\frac{A_{\mathrm{ph}}(\omega)}{W_{\mathrm{b}}},$ (31)
where we approximate the electron DOS $A(\varepsilon)$ with the inverse of the Hubbard bands’ bandwidth $W_{\mathrm{b}}$, introduced in Sec. 4. The strength of e-ph coupling is then encoded in the dimensionless parameter [83]
$\lambda_{\mathrm{eff}}=2\int_{0}^{\infty}\frac{d\omega}{\omega}\alpha^{2}F(\omega)=\frac{2g^{2}}{W_{\mathrm{b}}\omega_{\mathrm{ph}}},$ (32)
in which $\alpha^{2}F(\omega)$ from Eq. 31 depends now only on $\omega$. Finally, we notice that the Hartree shift and $\lambda$ are related by:
$\varepsilon_{\mathrm{H},z}=-\lambda_{\mathrm{eff}}W_{\mathrm{b}}\Delta n_{z}.$ (33)
@@ Appendix D Derivation of the time-averaged absorbed power
In this Appendix we derive the time-averaged absorbed power $\overline{P}_{\mathrm{abs}}$ introduced in Sec. 3.2 of the main text.
The time-dependent absorbed power $P_{\mathrm{abs}}(t)$ originates from the oscillating current $\bm{j_{\parallel}}(t)$ along the lattice body diagonal of the $d$-dimensional correlated layers, in the direction of the external electric field $\bm{E}(t)$.
We start from the expression of $P_{\mathrm{abs}}(t)$ as the classical expectation value
$P_{\mathrm{abs}}(t)=\frac{\mathop{}\!\mathrm{d}E_{\mathrm{abs}}}{\mathop{}\!\mathrm{d}t}=\Big{\langle}\frac{\mathop{}\!\mathrm{d}\hat{H}_{\mathrm{H}}(t)}{\mathop{}\!\mathrm{d}t}\Big{\rangle}.$ (34)
The time-dependence in $\hat{H}_{\mathrm{H}}(t)$  [^20] The subscript $\mathrm{H}$ stays for the Heisenberg representation. appears only via the vector potential $\bm{A}(t)$ in the intralayer hopping term $t_{z}(t)$ 2, introduced in Sec. 2. For this reason, one obtains the general result [61]
$P_{\mathrm{abs}}(t)=\Big{\langle}\frac{\partial\hat{H}_{\mathrm{H}}(t)}{\partial\bm{A}(t)}\cdot\frac{\partial\bm{A}(t)}{\partial t}\Big{\rangle}=\bm{j_{\parallel}}(t)\cdot\bm{E}(t),$ (35)
@@
$\bm{j}_{\parallel}(t)=-\Big{\langle}\frac{\partial\hat{H}_{\mathrm{H}}(t)}{\partial\bm{A}(t)}\Big{\rangle}=\frac{2\mathrm{i}}{N}\sum_{z=1}^{\mathrm{L}}\sum_{\bm{k}}\bm{v}_{\bm{k}}(t)G_{z,\bm{k}}^{<}(t,t),$ (36)
where $N$ is the number of lattice sites of the single $d$-dimensional correlated layer and $n_{z,\bm{k}}(t)=-\mathrm{i}G_{z,\bm{k}}^{<}(t,t)$.
@@
$\overline{P}_{\mathrm{abs}}=\frac{1}{T}\int_{-T/2}^{T/2}\mathop{}\!\mathrm{d}tP_{\mathrm{abs}}(t)$
$=\frac{2\mathrm{i}}{N}\sum_{z=1}^{\mathrm{L}}\sum_{\bm{k}}\frac{1}{T}\int_{-T/2}^{T/2}\mathop{}\!\mathrm{d}t\bm{v}_{\bm{k}}(t)\cdot\bm{E}(t)$
$\sum_{l=-\infty}^{\infty}e^{-\mathrm{i}l\Omega t}\int_{-\infty}^{+\infty}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}G_{l,z,\bm{k}}^{<}(\omega),$ (37)
@@
with $t_{\mathrm{rel}}=t-t^{\prime}$ and $t_{\mathrm{av}}=(t+t^{\prime})/2$ defined in Sec. 3.2. The scalar product in 37 reads
$\bm{v}_{\bm{k}}(t)\cdot\bm{E}(t)$ $=E_{0}\cos(\Omega t)\Big{\{}-\overline{\epsilon}\cos\Big{[}\mathscr{A}\sin(\Omega t)\Big{]}$ (39)
$+\epsilon\sin\Big{[}\mathscr{A}\sin(\Omega t)\Big{]}\Big{\}}$
with the quantities $\mathscr{A}$,$\epsilon$,$\overline{\epsilon}$ defined in Sec. 2. After some algebra, one finally obtains for the Wigner and layer resolved integrand in 37
$\overline{P}_{\mathrm{abs},l,z}(\omega)$ $=-E_{0}\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\rho(\epsilon,\overline{\epsilon})\Big{(}J_{1-l}(\mathscr{A})+J_{-1-l}(\mathscr{A})\Big{)}$ (40)
$\Big{[}\epsilon(1-\delta_{l,0})\delta_{l,2k}+\mathrm{i}\overline{\epsilon}\delta_{l,2k+1}\Big{]}G_{l,z}^{<}(\omega;\epsilon,\overline{\epsilon}),$
in which $k\in\mathbb{Z}$, $J_{n}$ is the $n$-th order Bessel function of the first kind and the integration over the JDOS $\rho(\epsilon,\overline{\epsilon})$ is used (see Sec. 2), which leads to Eq. 3.2 in Sec. 3.2.